# The Tidal Torque Theory Revisited. I. Protohalo Angular Momentum

### v2

@@ 3 Dominant Tidal Torque
The deformation tensor at the c.o.m. of the protohalo may be due to a mass excess identified to a peak or a mass defect identified to a hole, i.e. a local minimum. For simplicity we will concentrate for the moment on the former case, but we will come back to the latter possibility below.
From now on, the properties of any peak tracing the dominant peculiar mass excess causing a tidal torque on the protohalo are denoted with subindex ‘d’ to distinguish them from the properties of the peak associated with the protohalo. For instance, the mass of dominant peculiar mass excess with density contrast $\delta_{\rm d}$, different in general from $\delta$ (it is not necessarily collapsing at the same time as the protohalo), on the scale $R_{\rm d}$, different from $R$, is $M_{\rm d}=4\pi/3(R_{\rm d}^{\rm th})^{3}[1+\delta_{\rm d}(R_{\rm d})]\bar{\rho}(t_{\rm i})$.
There are, of course, many (actually infinite) nested peaks of different scales in the neighbourhood of the peak with $\delta$ at $R$ tracing the protohalo, and we cannot add all their tidal torques because they are included in that caused by one dominant object. But the CUSP formalism allows one to choose the relevant scale for each set of nested peaks. As shown by Manrique & Salvador-Solé (1995), nested peaks form continuous sets, the so-called peak trajectories, in the $R$-$\delta$ plane of the linear density field at $t_{\rm i}$, that trace the growth of different accreting haloes. [^2] Some peak trajectories are interrupted at some scale when the traced accreting halo suffers a major merger. But the fraction of interrupted peak trajectories is very small, so we may forget those cases in a first approximation. This means that all the peaks in every peak trajectory actually correspond to the same mass excess seen at different scales. We must thus concentrate on the scale at which each mass excess yields the strongest tidal torque at the c.o.m. of the protohalo, from now on the ’dominant scale’ of each peak.
A peak trajectory is given by the solution of the differential equation (Manrique & Salvador-Solé, 1995)
$\frac{{\rm d}\delta}{{\rm d}R}=-\langle x\rangle[\delta(R),R]\,\sigma_{2}(R)R,$ (21)
@@
where $\langle x\rangle(\delta,R)$ is the mean curvature of peaks with $\delta$ at $R$. For large $R$ as corresponding to the mass excesses under consideration, $\langle x\rangle$ is very approximately equal to $\gamma\nu$ (BBKS), so equation (22) becomes
$\frac{{\rm d}\ln\delta}{{\rm d}lnR}=m,$ (22)
where $m=-(n+3)/2$.
@@
Since the mass excess (or peculiar mass) of an ellipsoidal patch with density contrast $\delta_{\rm d}$ and scale $R_{\rm d}$ is $4\pi/3\,(R^{\rm th}_{\rm d})^{3}\delta_{\rm d}\bar{\rho}(t_{\rm i})$, to choose the peak along any given trajectory causing the strongest tidal torque we must find the scale $R_{\rm d}$ maximising $\delta(R_{\rm d})(R^{\rm th}_{\rm d})^{3}$ or $\delta(R_{\rm d})R_{\rm d}^{3}$, with $\delta(R_{\rm d})$ equal to the solution of equation (22)
$\delta(R_{\rm d})=\delta_{R}\,\left(\frac{R_{\rm d}}{R}\right)^{m}$ (23)
for the boundary condition $\delta_{\rm d}(R)=\delta_{R}$ for any given $\delta_{R}$, where $R$ is the scale of the peak marking the protohalo.
@@
From equation (23) we see that, for the CDM spectrum in the relevant mass range where $m\approx-0.63$, $\delta$ diminishes with increasing $R$ less rapidly than $R^{-3}$, so $\delta(R)R^{3}$ is ever increasing and has no maximum, actually. Thus, what really limits the scale $R_{\rm d}^{\rm th}\approx R_{\rm d}/2$ along any given peak trajectory is the separation $r$ to the c.o.m. of the protohalo. Indeed, for larger $R_{\rm d}$, the ellipsoidal shells lying beyond do not contribute to the tidal force at the c.o.m. of the protohalo, while $\delta$ is increasingly smaller, so the maximum tidal torque is found for the scale $R_{\rm d}^{\rm th}=r$. Note that, since $r>R^{\rm th}$, this implies that $R_{\rm d}>R$, so $\delta_{\rm d}$ is rigorously lower than $\delta$, implying that the protohalo will collapse before the dominant mass excess causing its tidal torque, which will still be in linear or mildly-non linear regime at that moment. This result will be used in Paper II.
Once the value of $R_{\rm d}$ is known, the peak trajectory (eq. [23]) provides the corresponding density contrast $\delta_{\rm d}=\delta(R_{\rm d})$. For simplicity in the notation, we omit from now on the explicit dependence of $\delta(R_{\rm d})$ on the boundary condition $\delta_{R}$ at $R$.
@@
The conditional PDF of peaks with $\nu_{\rm d}$ and $C_{\rm d}=\{e_{\rm d},p_{\rm d},\alpha_{\rm d},\beta_{\rm d},\kappa_{\rm d}\}$ properties in infinitesimal ranges at the dominant scale $R_{\rm d}$ lying at a distance $r$ from the peak with $\delta$ and $C=\{e,p,\alpha,\beta,\kappa\}$ at $R$, $P^{\rm d}(\nu_{\rm d},C_{\rm d},R_{\rm d},r|\delta,C,R)$, is
@@
where ${\cal N}_{\rm pk}(\nu_{\rm d},C_{\rm d},R_{\rm d}|\delta^{\prime},R)$ is the conditional number density of peaks with $\nu_{\rm d}=\delta_{\rm d}/\sigma_{0}(R_{\rm d})$ and $C_{\rm d}$ at $R_{\rm d}$, with $\delta_{\rm d}$ given by the trajectory solution of equation (23) at this scale, subject to lying on a point with $\delta^{\prime}$ at $R$, essentially equal to the point at the boundary $\delta_{R}=\delta^{\prime}$ at $R$ of that trajectory, and $P(\delta^{\prime},R|\delta,C,R,r)$ is the probability of finding such a point at $r$ from a peak with $\delta$ and $C$ at $R$. Note that the factor $1+\xi_{\rm pk}(r,\nu_{\rm d},R_{\rm d},\nu,R)]$, where $\xi_{\rm pk}(r;\delta_{\rm d},R_{\rm d},\delta,R)$ is the cross-correlation between peaks with $\delta$ at $R$ and with $\delta_{\rm d}$ at $R_{\rm d}$, enhances the local average conditional number density ${\cal N}_{\rm pk}(\nu_{\rm d},C_{\rm d},R_{\rm d}|\delta^{\prime},R)$ of peaks with $\delta_{\rm d}$ at $R_{\rm d}$ at a distance $r$ from a peak with $\delta$ at $R$ with respect to the global average conditional number density (i.e. with no constraint relative to the separation from any peak).
@@
where $P_{\rm E}(\alpha_{\rm d},\beta_{\rm d},\kappa_{\rm d})$ is the usual PDF of Euler angles and $P_{\rm ep}(e_{\rm d},p_{\rm d})$ is the ellipticity and prolateness-PDF of the form (12) for the peak marking the mass excess, but with $\langle x_{\rm d}\rangle$ replacing $x_{\rm d}$ and $G_{0}(\tilde{\gamma},\tilde{\gamma}\tilde{\nu})$ given by equation (14) for $i=0$ but with $\gamma$ and $\nu$ replaced by
@@
for some known function $A$ that is not relevant here, and variance [^3] The units of $r$ are different from those used in Manrique et al. (1998). Note also that, according to the normalisation used in the definition of $x$, the curvature variance $\langle[\Delta x]\rangle$ is unity.,
@@
where $\psi$ is the mass correlation function on scale $R$, $\xi(r)$, normalised to $\xi(0)$, $\psi^{\prime}$ is its $r$-derivative and $r$ is in units of $R_{\star}$. For high peaks, as corresponding to galactic haloes, $\langle x\rangle$ approaches $\gamma\nu$ (BBKS) and, for power-law spectra (with $n<-1$), the gradients of $\psi$ can be neglected in front of unity and of $\psi$, which in turn is of order unity ($r$ is small, of the order of $R$). Consequently, $\langle[\Delta\delta(r)]^{2}|C]\rangle$ approaches $\hat{\sigma}_{0}\equiv\sigma_{0}^{2}[1-\psi^{2}]\approx 0$ and $\langle\delta(r)|C\rangle\equiv\delta\psi\approx\delta$. Thus, the conditional probability $P(\delta^{\prime},R|\delta,C,R,r)$ approaches
@@
where $\delta_{\rm D}(\delta^{\prime}-\delta)$ is the Dirac delta function, does not depend neither on the properties $C$ nor on $r$ because they do not correlate with $\delta^{\prime}$. Thus, we can write $P(\delta^{\prime},R|\delta,C,R,r)\equiv P(\delta^{\prime},R|\delta,R)$.
With this expression for the conditional probability $P$, the integral on the right of equation (24) leads to
@@
where $\delta_{\rm d}$ and $R_{\rm d}$ are functions of $\delta$, $R$ (or $R^{\rm th}$) and $r$ through the relations $\delta_{\rm d}=(R_{\rm d}/R)^{m}\delta=(r/R^{\rm th})^{m}\delta$ (eqs. [8]-[9]). The second equality on the right of equation (31) arises from writing the number density of peaks with $\nu_{\rm d}$ at $R_{\rm d}$ that at $R$ have height $\nu$ directly in terms of the unconditional number density of peaks with $\nu$ at $R$ at the boundary of their trajectories. This in turn implies that the cross-correlation function of peaks with $\nu_{\rm d}$ at $R_{\rm d}$ and with $\nu$ at $R$ reduces to the autocorrelation function, $\xi_{\rm pk}(r)=B^{2}(R,\nu)\xi_{\rm m}(r)$, of peaks with $\nu$ at $R$. Given such a form of the right hand member of equation (31), $P^{\rm d}(\nu_{\rm d},C_{\rm d},R_{\rm d},r|\delta,R)$ is hereafter simply written as $P^{d}(C_{\rm d},r|\delta,R)$. Note the properties $C$ do not appear in the conditional PDF of the dominant mass excess properties, $P^{\rm d}$, though they will appear, of course, in the joint PDF of the ‘protohalo-main source of tidal torque’ properties used in Section 4.
Taking $\xi_{\rm pk}(r;\nu,R)$ given by the leading order of the perturbative bias expansion, $\xi_{\rm pk}(r;\nu,R)\approx B^{2}(R,\nu)\xi(r)$, where
$B(R,\nu)\approx 1.6\,\frac{\nu-\gamma\,G_{2}(\gamma,\gamma\nu)/G_{1}(\gamma,\gamma\nu)}{\sigma_{0}\,(1-\gamma^{2})}$ (32)
is the Lagrangian linear bias in the galactic halo mass range (Salvador-Solé & Manrique, 2024) and $\xi(r)=(r/s_{0})^{-\tilde{\gamma}}$, with $\tilde{\gamma}\approx 2$ and $s_{0}\sim 15h^{-1}$ Mpc (Abdullah et al., 2024), is the matter autocorrelation, the probability $P^{\rm d}$ takes the form
$P^{\rm d}(C_{\rm d}|\delta,R,r)=4\pi r^{2}\,[1+B^{2}(R,\nu)\xi(r)]{\cal N}_{\rm pk}(\nu,C_{\rm d},R),$ (33)
As mentioned, the previous derivation could also be followed for tidal torques caused by mass defects instead of mass excesses. Following the same derivation as in (31) for peaks but for holes (minima of the linear density field, i.e. like peaks but with negative $\nu$ and $x$), we would be led to a hole density contrast $\delta_{\rm h}=\tilde{B}\delta_{\rm m}$ with the linear bias of holes with $-\delta$ at $R$, $\tilde{B}$, equal to minus the linear bias of peaks with $\delta$ at $R$, $B$. This implies that, while the peak-peak correlation is given by $\xi_{\rm pk}(r)=B^{2}(R,\nu)\,\xi_{\rm m}(r)$, where $\xi_{\rm m}(r)$ is the matter correlation, the peak-hole cross-correlation is given by $\xi_{\rm pk\,h}(r)=-B^{2}(R,\nu),\xi_{\rm m}(r)$. Consequently, in the probability of finding the main source (peak or hole) of the tidal torque of the same modulus at a distance $r$ of the protohalo, $P^{\rm d}(C_{\rm d}|\delta,R,r)$ (eq. [33]), the term with the linear bias parameter $B^{2}$ arising from peaks cancels with that arising from holes, and we are led simply to
@@
Since the peculiar mass $M_{\rm d}$ of peaks or holes of the dominant scale along each trajectory with boundary condition $\delta$ or $-\delta$ at $R$ in the neighbourhood of the protohalo, is given by $4\pi/3\delta_{\rm d}(R_{\rm d})(R_{\rm d}^{\rm th})^{3}\bar{\rho}(t_{\rm i})$, with $R_{\rm d}^{\rm th}=r$, so the tidal force they produce on the protohalo is proportional to $r^{-3}$, their corresponding tidal torque is just proportional to their value $\delta_{\rm d}=\delta(R_{\rm d})$. And, since $\delta_{\rm d}$ decreases with increasing $R_{\rm d}\propto r$ according to equation (23) with $\delta_{R}=\delta$, the peak or hole causing the strongest tidal torque on the protohalo among all nearby peaks and holes of dominant scale is simply that lying at the smallest distance $r$ from it, i.e. with the smallest $R_{\rm d}$ and, hence, the highest $\delta_{\rm d}$. Consequently, we can concentrate from now in the nearest dominant peak or hole. Concretely, the probability that the nearest dominant peak or hole with $\nu_{\rm d}$ and $C_{\rm d}$ on scale $R_{\rm d}$ is at a distance $r$ from the c.o.m. of the protohalo is the probability $P^{\rm d}$ of finding one of such dominant peaks or holes at $r$ times the probability that there is no other dominant peak or hole closer to the c.o.m. of the protohalo,
@@
On the other hand, the peculiar gravitational potential $\Phi_{\rm d}$ of an ellipsoidal mass excess with uniform density $\delta_{\rm d}\bar{\rho}(t_{\rm i})$ at ${\bf x}=(x_{1},x_{2},x_{3})$ in physical Cartesian coordinates with origin at the c.o.m. of the protohalo and aligned with its principal axes is (e.g. Chandrasekhar 1987)
$\Phi_{\rm d}({\bf x})=\pi a_{\rm n1}a_{\rm n2}a_{\rm n3}G\delta_{\rm d}\bar{\rho}(t_{\rm i})~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$
×$\times\left[\int_{S({\bf x})}^{\infty}\frac{{\rm d}s}{\Delta(s)}-\sum_{i=1}^{3}x_{i}^{2}\int_{S({\bf x})}^{\infty}\frac{{\rm d}s}{(a_{{\rm n}i}^{2}+s)\Delta(s)}\right],~{}~{}~{}$ (36)
where $G$ is the gravitational constant, $a_{\rm n1}$, $a_{\rm n2}$ and $a_{\rm n3}$ are the top-hat semi-axes of the ellipsoidal mass excess, with total mass $M_{\rm d}=4\pi/3a_{\rm n1}a_{\rm n2}a_{\rm n3}\bar{\rho}(t_{\rm i})$,
$\Delta(s)\equiv\left[(a_{\rm n1}^{2}+s)(a_{\rm n2}^{2}+s)(a_{\rm n3}^{2}+s)\right]^{1/2}.$ (37)
@@
$\sum_{i=1}^{3}\frac{x_{i}^{2}}{a_{\rm ni}^{2}+S}=1.$ (38)
We remark that the expression (36) presumes that the point x is external to the ellipsoidal mass excess, which is indeed the case at the c.o.m. of the protohalo since, by construction, $R_{\rm d}^{\rm th}=r$. From that potential we can derive the Lagrangian tidal tensor acting on the protohalo at $r=R_{\rm d}^{\rm th}$ with all time-varying quantities scaled to the values at $t$, which is very approximately given by (see App. A)
${\bf T_{\rm d}}_{ij}=\frac{a^{3}(t_{\rm i})}{D(t_{\rm i})}\frac{4\pi}{3}G\delta_{\rm d}\bar{\rho}(t_{\rm i})\left[3\frac{x_{i}x_{j}}{a^{2}_{{\rm n}i}a^{2}_{{\rm n}j}}a_{\rm d}^{2}-\delta_{ij}\right],$ (39)
where $a_{\rm d}^{2}\equiv 1/3\sum_{k=1}^{3}a^{2}_{{\rm n}k}$. Note that in order to balance the dependence on $t_{\rm i}$ of $\delta_{\rm d}/r^{3}_{\rm d}$ and have a Lagrangian tidal tensor independent of $t_{\rm i}$, the tidal tensor calculated at that particular time has been multiplied by $a^{3}{t_{\rm i}}/D(t_{\rm i})$. The preceding derivation assumed a mass excess, but it also holds, of course, for the case of a mass defect (i.e. a virtual mass excess with negative $\delta_{\rm d}$).

### v5

@@ 3 Dominant Tidal Torque
The deformation tensor at the c.o.m. of the protohalo may be due to a mass excess identified to a peak or a mass default identified to a hole, i.e. a local minimum in density contrast, with identical statistial properties than peaks but for the opposite signs of the density contrast, curvature and eingenvalues of the secoond derivative tensor. For simplicity we will concentrate for the moment in the case of peaks, but we will come back to the most general case of peaks and holes at he end.
There are, of course, many (actually infinite) nested peaks of different scales in the neighbourhood of the peak with $\delta$ at $R$ tracing the protohalo and we cannot sum up all their individual tidal torques because the contribution of many of those peaks is included within that of a few others, that are dubbed here as the d͡ominant peak. From now on, the properties of those dominant peaks are denoted with subindex ‘d’ to distinguish them from the properties of the peak associated with the protohalo itself.
More specifically, as shown by Manrique & Salvador-Solé (1995), peaks are often nested within each other in continuous sets in the $R$-$\delta$ plane of the smoothed linear density field at $t_{\rm i}$, the so-called peak trajectories, tracing the mass growth of accreting haloes,  [^3] Some peak trajectories are interrupted at some scale when the corresponding accreting halo suffers a major merger. But such interruptions are little frequent, so we can ignore them in a first approximation. Peak trajectories satisfy the differential equation
$\frac{{\rm d}\delta^{\prime}}{{\rm d}R^{\prime}}=-\langle x\rangle[\delta(R^{\prime}),R^{\prime}]\,\sigma_{2}(R^{\prime})R^{\prime},$ (21)
@@
where $\langle x\rangle(\delta^{\prime},R^{\prime})$ is the mean curvature of peaks with $\delta^{\prime}$ at $R^{\prime}$. For large $R^{\prime}$ as corresponding to the mass excesses able to cause a substantial tidal torque to galaxy mass protohaloes, $\langle x\rangle$ is very approximately equal to $\gamma\nu$ (BBKS), so equation (22) becomes
$\frac{{\rm d}\ln\delta^{\prime}}{{\rm d}lnR^{\prime}}=m,$ (22)
@@
where $m=-(n+3)/2$. Since all peaks along a trajectory remain essentially at the same position (Manrique & Salvador-Solé, 1995) and the mass excess (peculiar mass) of the ellipsoidal patch with density contrast $\delta^{\prime}$ and scale $R^{\prime}$ is $4\pi/3\,[(R^{\rm th})^{\prime}]^{3}\delta^{\prime}\bar{\rho}(t_{\rm i})$, to choose the peak along any given trajectory causing the strongest tidal torque we must find the scale $R^{\prime}$ maximising $\delta^{\prime}(R^{\prime})[(R^{\rm th})^{\prime}]^{3}$ or $\delta^{\prime}(R^{\prime})(R^{\prime})^{3}$, with $\delta^{\prime}(R^{\prime})$ equal to the solution of equation (22)
$\delta^{\prime}(R^{\prime})=\delta_{R}\,\left(\frac{R^{\prime}}{R}\right)^{m}$ (23)
@@
for the boundary condition $\delta^{\prime}(R)=\delta_{R}$ with any given $\delta_{R}$, where $R$ is the scale of the peak marking the protohalo. From equation (23) we see that, for the CDM spectrum in the relevant mass range where $m\approx-0.63$, $\delta^{\prime}(R^{\prime})$ diminishes with increasing $R^{\prime}$ less rapidly than $(R^{\prime})^{-3}$, so $\delta(R^{\prime})[(R^{\prime})]^{3}$ increases monotonically with $R^{\prime}$. Therefore, there is no maximum $\delta^{\prime}(R^{\prime})(R^{\prime})^{3}$ value over a peak trajectory; what limits tidal torque is the scale $(R^{\rm th})^{\prime}\approx R^{\prime}/2$ which cannot exceed the separation $r$ to the c.o.m. of the protohalo. Indeed, for larger scales, the ellipsoidal shells lying beyond do not contribute to the tidal force at the c.o.m. of the protohalo (see below), while $\delta^{\prime}(R^{\prime})$ is increasingly smaller, so the maximum tidal torque setting the dominant peak is found for the scale $R^{\prime}=r$. That is thus the scale $R_{\rm d}$ of the dominant peak in each trajectory determined by the value of $\delta_{R}$ and its density contrast $\delta_{\rm d}$ is the corresponding value of $\delta^{\prime}(R_{\rm d})$. Note that, since $r>R^{\rm th}$ and, hence, $R_{\rm d}>R$, $\delta_{\rm d}\equiv\delta^{\prime}(R_{\rm d})$ is substantially smaller than the density contrast $\delta$ of the protohalo. Thus, when the protohalo will collapse, the dominant mass excess causing its tidal torque will still be in linear regime. This result will be used in Paper II.
@@
The conditional PDF of peaks with $\nu_{\rm d}$ and $C_{\rm d}=\{e_{\rm d},p_{\rm d},\alpha_{\rm d},\beta_{\rm d},\kappa_{\rm d}\}$ properties in infinitesimal ranges at the dominant scale $R_{\rm d}$ lying at a distance $r$ from the peak with $\delta$ and $C=\{e,p,\alpha,\beta,\kappa\}$ at $R$ is
@@
where ${\cal N}_{\rm pk}(\nu_{\rm d},C_{\rm d},R_{\rm d}|\delta^{\prime},R)$ is the conditional number density of peaks with $\nu_{\rm d}=\delta_{\rm d}/\sigma_{0}(R_{\rm d})$ and $C_{\rm d}$ at $R_{\rm d}$ subject to lying at a point with $\delta^{\prime}$ at $R$ or the point with $\delta^{\prime}=\delta_{R}$ at $R$ setting the boundary condition of the corresponding trajectory and $P(\delta^{\prime},R|\delta,C,R,r)$ is the probability of finding such a point at $r$ from a peak with $\delta$ and $C$ at $R$. The factor $1+\xi_{\rm pk}(r,\nu_{\rm d},R_{\rm d},\nu,R)]$, where $\xi_{\rm pk}(r;\delta_{\rm d},R_{\rm d},\delta,R)$ is the cross-correlation between peaks with $\delta$ at $R$ and with $\delta_{\rm d}$ at $R_{\rm d}$, enhances the local average conditional number density ${\cal N}_{\rm pk}(\nu_{\rm d},C_{\rm d},R_{\rm d}|\delta^{\prime},R)$ of peaks with $\delta_{\rm d}$ at $R_{\rm d}$ at a distance $r$ from a peak with $\delta$ at $R$ with respect to the global average conditional number density (i.e. with no constraint relative to the separation from any other peak).
@@
where $P_{\rm E}(\alpha_{\rm d},\beta_{\rm d},\kappa_{\rm d})$ is the usual PDF of Euler angles and $P_{\rm ep}(e_{\rm d},p_{\rm d})$ is the ellipticity- and prolateness-PDF of the form (12) for the dominant peak causing the tidal torque, but with $\langle x_{\rm d}\rangle$ replacing $x_{\rm d}$ for the above mentioned reason and $G_{0}(\tilde{\gamma},\tilde{\gamma}\tilde{\nu})$ given by equation (14) for $i=0$ but with $\gamma$ and $\nu$ replaced by
@@
for some known function $A$ that is not relevant here, and variance [^4] The units of $r$ are different from those used in Manrique et al. (1998). Note also that, according to the normalisation used in the definition of $x$, the curvature variance $\langle[\Delta x]\rangle$ is unity.
@@
where $\psi$ is the matter correlation function $\xi(r)$ on scale $R$ normalised to $\xi(0)$ and $\psi^{\prime}$ is its $r$-derivative, being $r$ in units of $R_{\star}$. For high peaks, as corresponding to galactic haloes, $\langle x\rangle$ approaches $\gamma\nu$ (BBKS) and, for power-law spectra (with $n<-1$), the gradients of $\psi$ can be neglected in front of unity and of $\psi$, which in turn is of order unity ($r$ is small, of the order of $R$). Consequently, $\langle[\Delta\delta(r)]^{2}|C]\rangle$ approaches $\hat{\sigma}_{0}\equiv\sigma_{0}^{2}[1-\psi^{2}]\approx 0$ and $\langle\delta(r)|C\rangle\equiv\delta\psi\approx\delta$. Thus, the conditional probability $P(\delta^{\prime},R|\delta,C,R,r)$ approaches
@@
where $\delta_{\rm D}(\delta^{\prime}-\delta)$ is the Dirac delta function and it does not depend neither on the properties $C$ nor on $r$ because they do not correlate with $\delta^{\prime}$. Thus, we can write $P(\delta^{\prime},R|\delta,C,R,r)\equiv P(\delta^{\prime},R|\delta,R)$.
With this approximate expression for the conditional probability $P$, the integral on the right of equation (24) leads to
@@
where $\delta_{\rm d}$ and $R_{\rm d}$ of the dominant peak are functions of $\delta$ and $R$ of the peak at $r$ from that marking the protohalo through the relations $\delta_{\rm d}=(R_{\rm d}/R)^{m}\delta=(r/R^{\rm th})^{m}\delta$ (eqs. [8]-[9]). The second equality on the right of equation (31) arises from writing the number density of peaks with $\nu_{\rm d}$ at $R_{\rm d}$ whose trajectory has a peak with $\delta$ at $R$ directly as the unconditional number density of peaks with $\nu$ at $R$. This in turn implies that the cross-correlation function of peaks with $\nu_{\rm d}$ at $R_{\rm d}$ and with $\nu$ at $R$ reduces to the autocorrelation function, $\xi_{\rm pk}(r)=B^{2}(R,\nu)\xi_{\rm m}(r)$, of peaks with $\nu$ at $R$. Given the form of the right hand member of equation (31), $P^{\rm d}(\nu_{\rm d},C_{\rm d},R_{\rm d},r|\delta,R)$ is simply $P^{d}(C_{\rm d},r|\delta,R)$. Note the properties $C$ do not appear in the conditional PDF of the dominant peak, $P^{\rm d}$, though they appear, of course, in the joint probability of the ‘protohalo-dominant tidal torque source’ properties used in Section 4.
Taking $\xi_{\rm pk}(r;\nu,R)$ given by the leading order of the perturbative bias expansion, $\xi_{\rm pk}(r;\nu,R)\approx B^{2}(R,\nu)\xi(r)$, where $B(R,\nu)\approx 1.6\,(\nu-\gamma\langle x\rangle)/[\sigma_{0}\,(1-\gamma^{2})]$ is protohalo linear bias in the galactic halo mass range (Salvador-Solé & Manrique, 2024), and $\xi(r)=(r/s_{0})^{-\tilde{\gamma}}$, with $\tilde{\gamma}\approx 2$ and $s_{0}\sim 15h^{-1}$ Mpc (Abdullah et al., 2024), is the matter autocorrelation, the probability $P^{\rm d}$ takes the form
$P^{\rm d}(C_{\rm d}|\delta,R,r)=4\pi r^{2}\,[1+B^{2}(R,\nu)\xi(r)]{\cal N}_{\rm pk}(\nu,C_{\rm d},R).$ (32)
As mentioned at the beginning of this Section, the previous derivation could also be followed for tidal torques caused by mass defaults instead of mass excesses. Following the same derivation for peaks leading to equation (31) but for holes (minima of the linear density field, i.e. like peaks but with negative $\nu$ and $x$), we would be led to a hole density contrast $\delta_{\rm h}=\tilde{B}\delta_{\rm m}$ with the linear bias of holes with $-\delta$ at $R$, $\tilde{B}$, equal to minus the linear bias of peaks with $\delta$ at $R$, $B$. This implies that, while the peak-peak correlation is given by $\xi_{\rm pk}(r)=B^{2}(R,\nu)\,\xi_{\rm m}(r)$, where $\xi_{\rm m}(r)$ is the matter correlation, the peak-hole cross-correlation is given by $\xi_{\rm pk\,h}(r)=-B^{2}(R,\nu),\xi_{\rm m}(r)$. Consequently, in the probability of finding the main source (peak or hole) of the tidal torque of the same modulus at a distance $r$ of the protohalo, $P^{\rm d}(C_{\rm d}|\delta,R,r)$ (eq. [32]), the term with the linear bias parameter $B^{2}$ arising from peaks cancels with that arising from holes, and we are led simply to
@@
The peculiar (excess of default) mass of the dominant peaks or holes along trajectories with boundary condition $|\delta(R)|$ in the neighbourhood of the protohalo is $M_{\rm d}=4\pi/3\delta_{\rm d}(R_{\rm d}^{\rm th})^{3}\delta_{\rm d}\bar{\rho}(t_{\rm i})$, with $R_{\rm d}^{\rm th}=r$. (For simplicity in the notation we skip the arguments of $\delta_{\rm d}(t_{\rm c},t_{\rm i}$), where $t_{\rm c}$ is the collapse time of the main tidal torque source.) Consequently, the corresponding tidal torque is just proportional to $\delta_{\rm d}=\delta^{\prime}(R_{\rm d})$. And, since $\delta_{\rm d}$ decreases with increasing $R_{\rm d}\propto r$ according to equation (23) with $\delta_{R}=\delta$, the main dominant peak or hole causing the strongest tidal torque on the protohalo is simply that lying the closest to it, i.e. with the smallest $R_{\rm d}$ and, hence, the highest $\delta_{\rm d}$. Thus, we will concentrate from now in the nearest dominant peak or hole, taking into account that the probability that the nearest dominant peak or hole with $\nu_{\rm d}$ and $C_{\rm d}$ on scale $R_{\rm d}$ is at a distance $r$ from the c.o.m. of the protohalo is the probability $P^{\rm d}$ of finding one of such dominant peaks or holes at $r$ times the probability that there is no other dominant peak or hole closer to the c.o.m. of the protohalo,
@@
On the other hand, the peculiar gravitational potential $\Phi_{\rm d}$ of the corresponding patch assumed with uniform peculiar density $\delta_{\rm d}\bar{\rho}(t_{\rm i})$ at ${\bf x}=(x_{1},x_{2},x_{3})$ in physical Cartesian coordinates with origin at the c.o.m. of the protohalo aligned with the principal axes of its ellipsoid is (e.g. Chandrasekhar 1987)
$\Phi_{\rm d}({\bf x})=\pi a_{\rm d1}a_{\rm d2}a_{\rm d3}G\delta_{\rm d}\bar{\rho}(t_{\rm i})~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$
×$\times\left[\int_{S({\bf x})}^{\infty}\frac{{\rm d}s}{\Delta(s)}-\sum_{i=1}^{3}x_{i}^{2}\int_{S({\bf x})}^{\infty}\frac{{\rm d}s}{(a_{{\rm d}i}^{2}+s)\Delta(s)}\right],~{}~{}~{}$ (35)
where we have written the peculiar (positive or negative) mass in terms of the semi-axes, $M_{\rm d}=4\pi/3a_{\rm d1}a_{\rm d2}a_{\rm d3}\delta_{\rm d}\bar{\rho}(t_{\rm i})$. In equation (35), $G$ is the gravitational constant, $a_{\rm d1}$, $a_{\rm d2}$ and $a_{\rm d3}$ are the top-hat semi-axes of the ellipsoidal mass excess,
$\Delta(s)\equiv\left[(a_{\rm d1}^{2}+s)(a_{\rm d2}^{2}+s)(a_{\rm d3}^{2}+s)\right]^{1/2}$ (36)
@@
$\sum_{i=1}^{3}\frac{x_{i}^{2}}{a_{{\rm d}i}^{2}+S}=1.$ (37)
Equation (35) presumes that the point x is external to the ellipsoidal mass excess or defect, which is indeed the case at the c.o.m. of the protohalo since, by construction, $R_{\rm d}^{\rm th}=r$. From that potential, we can derive the Lagrangian tidal tensor acting at $t$ on the protohalo at $r=R_{\rm d}^{\rm th}$ with all time-varying quantities scaled to the values at $t$, which is very approximately given by (see App. A)
${\bf T_{\rm d}}_{ij}=\frac{a^{3}(t_{\rm i})}{D(t_{\rm i})}\frac{4\pi}{3}G\delta_{\rm d}\bar{\rho}(t_{\rm i})\left[3\frac{x_{i}x_{j}}{a^{2}_{{\rm d}i}a^{2}_{{\rm d}j}}a_{\rm d}^{2}-\delta_{ij}\right],$ (38)
where $a_{\rm d}^{2}\equiv 1/3\sum_{k=1}^{3}a^{2}_{{\rm d}k}$. Without saying, the preceding derivation holds for positive (mass excesses) or negative (mass defaults) $\delta_{\rm d}$ values.